Use the References to access important values if needed for this question. Consider the reaction Fe(s) + 2 HCl(aq) FeCl2(s) + H2(g) Using the standard thermodynamic data in the tables linked above, calculate the equilibrium constant for this reaction at 298.15K. ANSWER:
Answer #1

Answer – We are given, reaction –

Fe(s) + 2 HCl(aq) -----> FeCl2(s) + H2(g)

First we need to calculate the ∆Gorxn

We know,

ΔGorxn = ∑ ΔGof of product – ∑ ΔGof of reactant

= [ (ΔGof FeCl2(s)) + (ΔGof H2(g) )] – [ ΔGof Fe (s) + 2*ΔGof HCl(aq)]

= (-302.34 + 0.00) – ( 0.00 + 2*-131.17)

= -42 kJ/mol

Now we know,

ΔGo = -RTlnK

-4.2*104 J = -8.314 J/mol.K * 298.15 K * ln K

So, ln K = -4.2*104 J / -8.314 J.mol-1 .K-1 * 298.15 K

              = 16.94

By taking antiln from both side

K = 2.28*107

The equilibrium constant for this reaction at 298.15K is 2.28*107